# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 3.1.1 May’s classical result
May considered that $\textbf{M}_{ij}$ are randomly drawn from a distribution with mean $\mu=0$ and variance $\sigma^{2}$ with probability $C$ and are 0 otherwise [163]. Hence, $\sigma$ represents the characteristic interspecies interaction strength and $C$ is the ratio between actual and potential interactions in the ecological system (often referred to as the connectance). For simplicity, the diagonal elements are chosen to be the same with $-d=-1$, representing the intrinsic damping timescale of each species, so that if disturbed from equilibrium, it would return with such a damping time by itself. May found that for random interactions drawn from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, a randomly assembled system is stable (in the sense that all the eigenvalues of the community matrix have negative real parts) if the so-called “complexity” measure
@@
This implies that more complexity (i.e., larger $CS$) tend to destabilize community dynamics [163, 165].
Table 3: Stability criteria for ecological systems with different interaction types. In each case, the stability criterion is derived for a large community matrix ${\bf M}$. Diagonal elements $\textbf{M}_{ii}$, representing self-regulation, are set to be $-d$. For random interactions, off-diagonal elements $\textbf{M}_{ij}$ are randomly drawn from a normal distribution $\mathcal{N}(0,\sigma^{2})$ with probability $C$ and are 0 otherwise. Predator-prey interactions come in pairs with opposite signs, i.e., $\textbf{M}_{ij}>0$ and $\textbf{M}_{ji}<0$. With probability $C$, we sample $\textbf{M}_{ij}$ from $|\mathcal{N}(0,\sigma^{2})|$ and $\textbf{M}_{ji}$ from $-|\mathcal{N}(0,\sigma^{2})|$. With probability $(1-C)$, both $\textbf{M}_{ij}$ and $\textbf{M}_{ji}$ are set to be 0. Community matrices with mutualistic and/or competitive interactions can be constructed similarly [11].
@@
May’s result is closely related to Girko’s Circular Law [104] in random matrix theory. Consider a random $S\times S$ real matrix ${\bf M}$ with entries independent and taken randomly from a normal distribution $\mathcal{N}(0,\sigma^{2})$. Then as $S\to\infty$, the eigenvalues of ${\bf M}/\sqrt{S\sigma^{2}}$ are uniformly distributed in the unit disk centered at $(0,0)$ in the complex plane [103]. Sommers et al. considered possible correlations between the off-diagonal elements $\textbf{M}_{ij}$ and $\textbf{M}_{ji}$ and proved that the eigenvalues of M are uniformly distributed in an ellipse with the real and imaginary directions $1+\rho$ and $1-\rho$ (where $\rho=\mathbb{E}[\textbf{M}_{ij}\textbf{M}_{ji}]/\text{Var}(\textbf{M}_{ij})$ for $i\neq j$), respectively, known as Sommers’ Elliptical Law [239].
Figure 3: Eigenvalue distributions of 10 community matrices (color) with $-d=-1$ on the diagonal and off-diagonal elements following the random (a), predator-prey (b), and mixture of competition and mutualism interactions (c), respectively. $S=250$, $C=0.25$, $\sigma=1$. The black ellipses are analytical results. Panels a, b, and c were redrawn from [11]. Impact of degree heterogeneity on the stability of ecological systems with random interactions (d), predator-prey (e), and mixture of competition and mutualism interactions (f). The dots represent the results from numerical simulations on 3-modal networks with different connectances. Higher $\text{Re}[\lambda_{1}]$ indicates lower stability. Each error bar represents the standard deviation of 100 independent runs. $S=1200$, $\sigma=1$, and $d=0$. Panel d was drawn in the log-log scale. Panels d, e, and f were redrawn from [270].
May’s result continues to be influential almost five decades later, not because it asserts that ecological systems must be inherently unstable, but rather because it highlights the importance of specific structural characteristics that enable real ecological systems to maintain stability despite their inherent complexity [39]. In other words, nature must adopt some devious and delicate strategies to cope with this stability-complexity paradox.
@@ 3.1.2 Impacts of interaction types and correlation
One of the specificity is the existence of well-defined interspecific relationships observed in nature, e.g., predator-prey, competition, mutualism, and a mixture of competition and mutualism. In 2012, by leveraging Sommers’ Ellipical Law [239], Allesina et al. refined May’s result and provided stability criteria for all these interspecific interaction types [11] , as shown in Table 3 and Fig. 3a, b, and c. They found remarkable differences between predator-prey interactions, which are stabilizing, but mutualistic and competitive interactions, which are destabilizing. Additionally, the correlation between interaction strengths determines interaction types. Tang et al. therefore incorporated correlation into the stability analysis by deriving a new stability criterion for large ecological systems with random interactions [249] , i.e.,
@@
where $\rho$ and $\mu$ are the overall correlation between pairs of interactions and the mean of the off-diagonal elements in the community matrix M, respectively. The criterion (5) can be viewed as a generalization of those for predator-prey, competition, and mutualism, derived by Allesina et al. [11]. The authors found that the effect of correlation of interaction strengths substantially influences the stability of large food webs (with predator-prey interactions) compared to other network structural properties. Hence, the presence of correlation between interactions of species can significantly influence the locations of the eigenvalues of the community matrix and the resulting stability of ecological systems [11, 37, 147, 249].
@@ 3.1.3 Impacts of degree heterogeneity
Many of the “devious strategies” adopted by nature can now be tested with the revised formula as a reference point. For example, one can further study the impact of degree heterogeneity on the stability of ecological systems. Degree heterogeneity measures the variability of the number of interactions associated with each species. Yan et al. found that for ecological systems with random interactions or a mixture of competition and mutualism interactions, increasing the degree of heterogeneity always destabilizes ecological systems [270], see Fig. 3d, f. For ecological systems with predator-prey interactions constructed from either simple network models or a realistic food web model (cascade model), high heterogeneity is always destabilizing, yet moderate heterogeneity is stabilizing [269, 270], see Fig. 3e. Furthermore, they obtained a stability criterion for large ecological systems with random interactions by considering the factor of degree heterogeneity, i.e.,
@@
where $\xi$ denotes the degree heterogeneity. Similarly, Allesina et al. found that broad degree distributions tend to stabilize food webs by approximating the real part of the leading (“rightmost”) eigenvalue of the community matrix [9]. Recently, Baron derived a closed-form expression for the eigenvalue spectrum of a general directed and weighted network [35]. The findings of this study suggest that network heterogeneity appears to be a destabilizing influence in most circumstances, except when the interactions are very asymmetric (e.g., very asymmetric predator-prey interactions). These results are consistent with what were reported by Yan et al. [270].
@@ 3.1.4 Impacts of self-regulation
for $i=1,2,\dots,S$, where $\lambda_{1}$ is the leading (“rightmost”) eigenvalue of the community matrix M without self-regulations [30]. Notably, empirical food webs can only achieve stability when the majority of species exhibit strong self-regulation [30]. Even for random ecological systems, attaining stability also requires negative self-regulations for a large amount of species [30]. In addition, Tang et al. investigated the effect of non-constant diagonal on the eigenvalue distribution of the community matrix M [249]. They found that a moderate variance of the diagonal elements minimally affects the distribution of the eigenvalues. Therefore, if the diagonal elements of M follow a moderate-variance distribution with mean satisfying (7), the ecological system will remain locally stable. Nevertheless, when the variance significantly surpasses that of the interspecific interactions, the impact on stability depends on the specific patterns of the diagonal elements. When the self-regulation is stronger for species with fewer interactions, the impact of a considerable variance on stability remains negligible [249].
@@ 3.1.5 Impacts of modularity
The stability of ecological systems can also be influenced by network modularity. A modular network can be divided into different modules or subsystems, and the interactions within each subsystem are much more frequent than those between subsystems [178]. The modularity of a network can be defined as
@@
where $L_{w}$ is the observed number of interactions within the subsystems, and $L_{b}$ is the number of inter-subsystem interactions. Grilli et al. studied the effect of modularity $Q$ on the stability of ecological systems [112]. They found that modularity exhibits a moderate stabilizing effect when the subsystems have similar sizes and the overall mean interaction strength is negative. In particular, the stabilizing effect becomes stronger for negative correlations. Conversely, anti-modularity is highly destabilizing, except for the case where the overall mean interaction strength is close to zero. The authors further investigated the effect of modularity in food webs through numerical simulations and discovered that the results remain qualitatively unchanged [112].
@@ 3.1.6 Impacts of dispersal
Spatial flows (e.g., exchanges of individuals, energy, and material) among local ecosystems are ubiquitous in nature [204]. Gravel et al. incorporated dispersal, i.e., spatial movement of species among local ecological systems, into the stability analysis of meta-ecological systems [109]. The community matrix M of a meta-ecological system can be expressed as the sum of three matrices, i.e.,
@@
where D is a diagonal matrix that accounts for intraspecific density dependence, Q is a matrix representing dispersal among patches with diffusion coefficient $q$, and K is a block diagonal matrix that contains the local community matrices (random matrices with $S$ species, connectance $C$, and interspecific interaction strength $\sigma$). By assuming the number of local systems and $S$ are large, the authors obtained the following stability criterion [109]:
@@
Therefore, the effect of dispersal can promote stability in meta-ecological systems, in which dispersal can move the most of the eigenvalues of the community matrix towards more negative values and shrinks the range of the remaining in proportion to the number of effective patches. Interestingly, Baron et al. discovered that the introduction of dispersal can lead to the Turing-type instability (i.e., the equilibrium point is unstable with respect to disturbances of a finite range of wavelengths) in an ecological system with a trophic structure [36]. While the inclusion of trophic structures often enhances stability in large ecological systems (e.g., food webs) [114, 133], it can lead to a peak in the real part of the maximum eigenvalue of the community matrix in the presence of dispersal, rendering the equilibrium point unstable [36].
@@ 3.1.7 Impacts of time delay
Empirical evidence has indicated that species interactions often exhibit time lags, rather than occurring instantaneously [179, 182]. Thus, time delays can hold significant implications for stability and coexistence. Notably, Pigani et al. investigated delay effects on the stability of large ecological systems [198], which can be captured by the following linearized equation:
@@ 3.2 Model-explicit approaches
The most commonly used model to describe the dynamics of ecological systems is the generalized Lotka-Volterra (GLV) model [49, 160] defined as
@@
Figure 4: Eigenvalue distributions of the three matrices M, Q, and J. The off-diagonal elements of A are sampled from a normal bivariate distribution with identical marginal $\mu_{A}=5/S$, $\sigma_{A}=5/{\sqrt{S}}$, and correlation $\rho_{A}=-0.5$. The diagonal elements of A are fixed to $\mu_{d}=-1$, while the diagonal elements of X are sampled from a uniform distribution on [0,1]. The two matrices Q and J are defined as $\textbf{Q}=\textbf{X}[(\mu_{d}-\mu_{A})\textbf{I}+\mu_{A}\textbf{1}]$ and $\textbf{J}=\textbf{X}(\mu_{d}\textbf{I}+\textbf{B})$, respectively, such that the bulk of the eigenvalue distributions of M are the same as these of J, and the outlier eigenvalue of M is the same as that of Q. Here, B is a random matrix with zero diagonal elements whose off-diagonal elements have mean zero and variance $\sigma_{A}^{2}$. $S=500$. This figure was redrawn from [100].
@@ 3.2.1 Impacts of equilibria
Equilibrium points play a significant role in the local stability analysis of the GLV model, as its community matrix is directly related to the equilibrium points of the system. Gibbs et al. explored the eigenvalue spectrum of the community matrix by assuming that the diagonal equilibrium matrix X and the interaction matrix A are randomly drawn from an arbitrary distribution with positive support and a bivariate distribution, respectively [100]. They proved that the eigenvalue spectrum of the community matrix M consists of a bulk of eigenvalues with mean $\mu_{X}(\mu_{d}-\mu_{A})$ determined by the eigenvalues of the matrix $\textbf{J}=\textbf{X}(\mu_{d}\textbf{I}+\textbf{B})$ and an outlier eigenvalue determined by the largest eigenvalue of the matrix $\textbf{Q}=\textbf{X}[(\mu_{d}-\mu_{A})\textbf{I}+\mu_{A}\textbf{1}]$ (1 is the all-one matrix) such that
@@
where $\mu_{X}$, $\mu_{d}$, and $\mu_{A}$ represent the mean of the diagonal of X, the diagonal of A, and the off-diagonal of A, respectively, see Fig. 4. Based on their finding, the authors concluded that for mutualistic ecological systems, if the interaction matrix is stable, then the community matrix will also be stable [100]. It is important to note that Gibbs et al. made the assumption that the distribution from which the equilibrium is drawn is independent of the interaction matrix, which does not typically hold in reality. Nevertheless, Liu et al. discovered that if the equilibrium point follows a specific distribution related to the elements of the interaction matrix A, Gibbs et al.’s assumption remains valid [155]. Around the same time, Stone demonstrated that for large ecological systems described by the GLV model, the stability of the interaction matrix implies the stability of the community matrix, under the condition that all species have positive equilibria [244].

### 07 Feb 2024 v2

@@ 3.1.1 May’s classical result
May considered that $\textbf{M}_{ij}$ are randomly drawn from a distribution with mean $\mu=0$ and variance $\sigma^{2}$ with probability $C$ and are 0 otherwise May-Nature-1972 . Hence, $\sigma$ represents the characteristic interspecies interaction strength and $C$ is the ratio between actual and potential interactions in the ecological system (often referred to as the connectance). For simplicity, the diagonal elements are chosen to be the same with $-d=-1$, representing the intrinsic damping timescale of each species, so that if disturbed from equilibrium, it would return with such a damping time by itself. May found that for random interactions drawn from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, a randomly assembled system is stable (in the sense that all the eigenvalues of the community matrix have negative real parts) if the so-called “complexity” measure
@@
This implies that more complexity (i.e., larger $CS$) tend to destabilize community dynamics May-Nature-1972 ; May-Book-1974 .
Table 3: Stability criteria for ecological systems with different interaction types. In each case, the stability criterion is derived for a large community matrix ${\bf M}$. Diagonal elements $\textbf{M}_{ii}$, representing self-regulation, are set to be $-d$. For random interactions, off-diagonal elements $\textbf{M}_{ij}$ are randomly drawn from a normal distribution $\mathcal{N}(0,\sigma^{2})$ with probability $C$ and are 0 otherwise. Predator-prey interactions come in pairs with opposite signs, i.e., $\textbf{M}_{ij}>0$ and $\textbf{M}_{ji}<0$. With probability $C$, we sample $\textbf{M}_{ij}$ from $|\mathcal{N}(0,\sigma^{2})|$ and $\textbf{M}_{ji}$ from $-|\mathcal{N}(0,\sigma^{2})|$. With probability $(1-C)$, both $\textbf{M}_{ij}$ and $\textbf{M}_{ji}$ are set to be 0. Community matrices with mutualistic and/or competitive interactions can be constructed similarly Allesina-Nature-2012 .
@@
May’s result is closely related to Girko’s Circular Law girko2004strong in random matrix theory. Consider a random $S\times S$ real matrix ${\bf M}$ with entries independent and taken randomly from a normal distribution $\mathcal{N}(0,\sigma^{2})$. Then as $S\to\infty$, the eigenvalues of ${\bf M}/\sqrt{S\sigma^{2}}$ are uniformly distributed in the unit disk centered at $(0,0)$ in the complex plane Girko-TPA-1984 . Sommers et al. considered possible correlations between the off-diagonal elements $\textbf{M}_{ij}$ and $\textbf{M}_{ji}$ and proved that the eigenvalues of M are uniformly distributed in an ellipse with the real and imaginary directions $1+\rho$ and $1-\rho$ (where $\rho=\mathbb{E}[\textbf{M}_{ij}\textbf{M}_{ji}]/\text{Var}(\textbf{M}_{ij})$ for $i\neq j$), respectively, known as Sommers’ Elliptical Law Sommers-PRL-1988 .
Figure 3: Eigenvalue distributions of 10 community matrices (color) with $-d=-1$ on the diagonal and off-diagonal elements following the random (a), predator-prey (b), and mixture of competition and mutualism interactions (c), respectively. $S=250$, $C=0.25$, $\sigma=1$. The black ellipses are analytical results. Panels a, b, and c were redrawn from Allesina-Nature-2012 . Impact of degree heterogeneity on the stability of ecological systems with random interactions (d), predator-prey (e), and mixture of competition and mutualism interactions (f). The dots represent the results from numerical simulations on 3-modal networks with different connectances. Higher $\text{Re}[\lambda_{1}]$ indicates lower stability. Each error bar represents the standard deviation of 100 independent runs. $S=1200$, $\sigma=1$, and $d=0$. Panel d was drawn in the log-log scale. Panels d, e, and f were redrawn from yan2017degree .
May’s result continues to be influential almost five decades later, not because it asserts that ecological systems must be inherently unstable, but rather because it highlights the importance of specific structural characteristics that enable real ecological systems to maintain stability despite their inherent complexity Bascompte-Science-2010 . In other words, nature must adopt some devious and delicate strategies to cope with this stability-complexity paradox.
@@ 3.1.2 Impacts of interaction types and correlation
One of the specificity is the existence of well-defined interspecific relationships observed in nature, e.g., predator-prey, competition, mutualism, and a mixture of competition and mutualism. In 2012, by leveraging Sommers’ Ellipical Law Sommers-PRL-1988 , Allesina et al. refined May’s result and provided stability criteria for all these interspecific interaction types Allesina-Nature-2012 , as shown in Table 3 and Figure 3a, b, and c. They found remarkable differences between predator-prey interactions, which are stabilizing, but mutualistic and competitive interactions, which are destabilizing. Additionally, the correlation between interaction strengths determines interaction types. Tang et al. therefore incorporated correlation into the stability analysis by deriving a new stability criterion for large ecological systems with random interactions tang2014correlation , i.e.,
@@
where $\rho$ and $\mu$ are the overall correlation between pairs of interactions and the mean of the off-diagonal elements in the community matrix M, respectively. The criterion (5) can be viewed as a generalization of those for predator-prey, competition, and mutualism, derived by Allesina et al. Allesina-Nature-2012 . The authors found that the effect of correlation of interaction strengths substantially influences the stability of large food webs (with predator-prey interactions) compared to other network structural properties. Hence, the presence of correlation between interactions of species can significantly influence the locations of the eigenvalues of the community matrix and the resulting stability of ecological systems Allesina-Nature-2012 ; baron2022eigenvalues ; kuczala2016eigenvalue ; tang2014correlation .
@@ 3.1.3 Impacts of degree heterogeneity
Many of the “devious strategies” adopted by nature can now be tested with the revised formula as a reference point. For example, one can further study the impact of degree heterogeneity on the stability of ecological systems. Degree heterogeneity measures the variability of the number of interactions associated with each species. Yan et al. found that for ecological systems with random interactions or a mixture of competition and mutualism interactions, increasing the degree of heterogeneity always destabilizes ecological systems yan2017degree , see Figure 3d, f. For ecological systems with predator-prey interactions constructed from either simple network models or a realistic food web model (cascade model), high heterogeneity is always destabilizing, yet moderate heterogeneity is stabilizing Yan-arXiv-14 ; yan2017degree , see Figure 3e. Furthermore, they obtained a stability criterion for large ecological systems with random interactions by considering the factor of degree heterogeneity, i.e.,
@@
where $\xi$ denotes the degree heterogeneity. Similarly, Allesina et al. found that broad degree distributions tend to stabilize food webs by approximating the real part of the leading (“rightmost”) eigenvalue of the community matrix allesina2015predicting . Recently, Baron derived a closed-form expression for the eigenvalue spectrum of a general directed and weighted network baron2022eigenvalue . The findings of this study suggest that network heterogeneity appears to be a destabilizing influence in most circumstances, except when the interactions are very asymmetric (e.g., very asymmetric predator-prey interactions). These results are consistent with what were reported by Yan et al. yan2017degree .
@@ 3.1.4 Impacts of self-regulation
for $i=1,2,\dots,S$, where $\lambda_{1}$ is the leading (“rightmost”) eigenvalue of the community matrix M without self-regulations barabas2017self . Notably, empirical food webs can only achieve stability when the majority of species exhibit strong self-regulation barabas2017self . Even for random ecological systems, attaining stability also requires negative self-regulations for a large amount of species barabas2017self . In addition, Tang et al. investigated the effect of non-constant diagonal on the eigenvalue distribution of the community matrix M tang2014correlation . They found that a moderate variance of the diagonal elements minimally affects the distribution of the eigenvalues. Therefore, if the diagonal elements of M follow a moderate-variance distribution with mean satisfying (7), the ecological system will remain locally stable. Nevertheless, when the variance significantly surpasses that of the interspecific interactions, the impact on stability depends on the specific patterns of the diagonal elements. When the self-regulation is stronger for species with fewer interactions, the impact of a considerable variance on stability remains negligible tang2014correlation .
@@ 3.1.5 Impacts of modularity
The stability of ecological systems can also be influenced by network modularity. A modular network can be divided into different modules or subsystems, and the interactions within each subsystem are much more frequent than those between subsystems newman2006modularity . The modularity of a network can be defined as
@@
where $L_{w}$ is the observed number of interactions within the subsystems, and $L_{b}$ is the number of inter-subsystem interactions. Grilli et al. studied the effect of modularity $Q$ on the stability of ecological systems grilli2016modularity . They found that modularity exhibits a moderate stabilizing effect when the subsystems have similar sizes and the overall mean interaction strength is negative. In particular, the stabilizing effect becomes stronger for negative correlations. Conversely, anti-modularity is highly destabilizing, except for the case where the overall mean interaction strength is close to zero. The authors further investigated the effect of modularity in food webs through numerical simulations and found that the results remain qualitatively unchanged grilli2016modularity .
@@ 3.1.6 Impacts of dispersal
Spatial flows (e.g., exchanges of individuals, energy, and material) among local ecosystems are ubiquitous in nature polis1995extraordinarily . Gravel et al. incorporated dispersal, i.e., spatial movement of species among local ecological systems, into the stability analysis of meta-ecological systems gravel2016stability . The community matrix M of a meta-ecological system can be expressed as the sum of three matrices, i.e.,
@@
where D is a diagonal matrix that accounts for intraspecific density dependence, Q is a matrix representing dispersal among patches with diffusion coefficient $q$, and K is a block diagonal matrix that contains the local community matrices (random matrices with $S$ species, connectance $C$, and interspecific interaction strength $\sigma$). By assuming the number of local systems and $S$ are large, the authors obtained the following stability criterion gravel2016stability :
@@
Therefore, the effect of dispersal can promote stability in meta-ecological systems, in which dispersal can move the most of the eigenvalues of the community matrix towards more negative values and shrinks the range of the remaining in proportion to the number of effective patches. Interestingly, Baron et al. discovered that the introduction of dispersal can lead to the Turing-type instability (i.e., the equilibrium point is unstable with respect to disturbances of a finite range of wavelengths) in an ecological system with a trophic structure baron2020dispersal . While the inclusion of trophic structures often enhances stability in large ecological systems (e.g., food webs) gross2009generalized ; johnson2014trophic , it can lead to a peak in the real part of the maximum eigenvalue of the community matrix in the presence of dispersal, rendering the equilibrium point unstable baron2020dispersal .
@@ 3.1.7 Impacts of time delay
Empirical evidence has indicated that species interactions often exhibit time lags, rather than occurring instantaneously niebur1991collective ; ohira1999resonance . Thus, time delays can hold significant implications for stability and coexistence. Notably, Pigani et al. investigated delay effects on the stability of large ecological systems pigani2022delay , which can be captured by the following linearized equation:
@@ 3.2 Model-explicit approaches
The most commonly used model to describe the dynamics of ecological systems is the generalized Lotka-Volterra (GLV) model bhargava1989generalized ; malcai2002theoretical defined as
@@
Figure 4: Eigenvalue distributions of the three matrices M, Q, and J. The off-diagonal elements of A are sampled from a normal bivariate distribution with identical marginal $\mu_{A}=5/S$, $\sigma_{A}=5/{\sqrt{S}}$, and correlation $\rho_{A}=-0.5$. The diagonal elements of A are fixed to $\mu_{d}=-1$, while the diagonal elements of X are sampled from a uniform distribution on [0,1]. The two matrices Q and J are defined as $\textbf{Q}=\textbf{X}[(\mu_{d}-\mu_{A})\textbf{I}+\mu_{A}\textbf{1}]$ and $\textbf{J}=\textbf{X}(\mu_{d}\textbf{I}+\textbf{B})$, respectively, such that the bulk of the eigenvalue distributions of M are the same as these of J, and the outlier eigenvalue of M is the same as that of Q. Here, B is a random matrix with zero diagonal elements whose off-diagonal elements have mean zero and variance $\sigma_{A}^{2}$. $S=500$. This figure was redrawn from gibbs2018effect .
@@ 3.2.1 Impacts of equilibria
Equilibrium points play a significant role in the local stability analysis of the GLV model, as its community matrix is directly related to the equilibrium points of the system. Gibbs et al. explored the eigenvalue spectrum of the community matrix by assuming that the diagonal equilibrium matrix X and the interaction matrix A are randomly drawn from an arbitrary distribution with positive support and a bivariate distribution, respectively gibbs2018effect . They proved that the eigenvalue spectrum of the community matrix M consists of a bulk of eigenvalues with mean $\mu_{X}(\mu_{d}-\mu_{A})$ determined by the eigenvalues of the matrix $\textbf{J}=\textbf{X}(\mu_{d}\textbf{I}+\textbf{B})$ and an outlier eigenvalue determined by the largest eigenvalue of the matrix $\textbf{Q}=\textbf{X}[(\mu_{d}-\mu_{A})\textbf{I}+\mu_{A}\textbf{1}]$ (1 is the all-one matrix) such that
@@
where $\mu_{X}$, $\mu_{d}$, and $\mu_{A}$ represent the mean of the diagonal of X, the diagonal of A, and the off-diagonal of A, respectively, see Figure 4. Based on their finding, the authors concluded that for mutualistic ecological systems, if the interaction matrix is stable, then the community matrix will also be stable gibbs2018effect . It is important to note that Gibbs et al. made the assumption that the distribution from which the equilibrium is drawn is independent of the interaction matrix, which does not typically hold in reality. Nevertheless, Liu et al. discovered that if the equilibrium point follows a specific distribution related to the elements of the interaction matrix A, Gibbs et al.’s assumption remains valid liu2023feasibility . Around the same time, Stone demonstrated that for large ecological systems described by the GLV model, the stability of the interaction matrix implies the stability of the community matrix, under the condition that all species have positive equilibria stone2018feasibility .